1. 100 kmols of moist hydrogen gas containing 4% water by mole is burnt completely in a furnace with 32% excess dry air. a) Complete analysis
b) Orsat analysis
2. A furnace is charged with 100 kmols of fuel gas, the composition of which is: 50% CO, 30% H2, 5% CH4, 5% C2H6 and 10% O2. Dry air is supplied 50% in excess for combustion. Assuming complete combustion.
a) Complete analysis
b) Orsat analysis
1
Expert's answer
2020-12-01T09:28:22-0500

1. 2H2 + O2 = 2H2O

100 x 0.96 = 96 kmols H2.

96/20 x 0.68= 70,5882353 kmols O2.

4 kmols H2O.

2. 2CO + O2 = 2CO2

2H2 + O2 = 2H2O

CH4 + 2O2 = CO2 + 2H2O

C2H6 + 2.5O2 = CO2 + 3H2O

50/2= 25 O2. 50 CO2.

30 /2= 15 O2. 30 H2O.

5x2 = 10 O2. 5 CO2 and 10 H2O.

5x2 = 10 O2. 5 CO2 and 10 H2O.

25+15+10+10=60 kmols O2

60/0.5 = 120. 120 + 6 = 126 O2.